Suppose you wish to fabricate a uniform wire from 2.30 g of copper. The wire is to have a resistance of

R = 0.580 Ω and all the copper is to be used.

(a) What must be the length of this wire?
2.96m
(b) What must be the diameter of this wire?
Answer #1

(a) Density of copper is ρ = 8.94 grams/cm³ and its resistivity σ = 16.78 * 10^(-9) Ω·m

V = m/ρ

AL = m/ρ

A = m/(ρL)

R = σL/A

R = σL²ρ/m

0.580 = 16.78 * 10^(-9) L²(8.94 * 1000)/(2.3/1000)

L = 2.98 m

(b).

AL = m/ρ

(¼ π d²)L = m/ρ

(¼ π d²)(2.98) = (2.30 * 10^(-3))/(8.94 * 1000)

d = 3.31*10^-4 m

(or d = 0.331 mm

or 331 um)

Kindly upvote :)